# LuSEE-Night Science Collaboration

### v1

@@ Abstract
preprint: arXiv:xxxx.xxxx
@@ I Introduction
The lunar far side has been touted as the ultimate radio observatory for low-frequency (<30<30 MHz) observations of the sky. At these low frequencies, observations from Earth are limited by both the Earth’s ionosphere, which refracts and reflects radiation from the sky, and emissions from anthropogenic and meteorological sources. The lunar far side, permanently shielded from Earth’s signals and, during the lunar night, also shielded from the Sun, has long been considered an ideal location for low-frequency radio receivers. Indeed, there are several proposals for large futuristic telescopes, such as FarSide [1], ALO-DEX [2], LCRT [3], as well as several recent operating pathfinders, such as the receiver on Chang’e-4 [4], ROLSES [5, 6] on the Intuitive Machines 1 mission, and the upcoming LuSEE-Night on the Firefly’s Blue Ghost 2 mission.
LuSEE-Night is a pathfinder radio telescope slated for delivery to the lunar far side in September 2026. A preliminary instrument description can be found in [7], with a more complete instrument design paper in preparation. Since the monopole antennas are electrically short at the lower end of the observed frequency range and have inherently low directivity, the instrument is more akin to a radio receiver than a traditional telescope. Given the available 16 correlation products, coupled with lunar rotation and rotation of the antenna turntable, however, the dataset will contain a large number of distinct measurements. Although each one of these measurements probes large patches of the sky, their number, and signal-to-noise will overconstrain the degrees of freedom available in the low-resolution map of the sky. The purpose of this paper is to investigate this problem quantitatively and to determine if LuSEE-Night’s data can be used to recover a map of the radio sky.
To this end, we employ the Wiener filter, a well-established linear method for map reconstruction. Wiener filtering provides the maximum a posteriori (MAP) map estimate assuming a Gaussian distribution for the input map pixels and noise. It has been traditionally used in Cosmic Microwave Background (CMB) map-making [8] and for general data compression in cosmology (e.g., [9]). In this work, we utilize Wiener filtering as a deconvolution process under a minimal set of assumptions. Future work (e.g., [10]) will explore more aggressive non-linear map-making algorithms that may produce higher fidelity maps, albeit at the cost of stronger, and potentially incorrect, assumptions.
@@ II.4 Simulating the data vector
For the input sky model, we use the Ultra-Low frequency Sky Model (ULSA) [11], which we regenerate at frequencies from 1 to 50 MHz with a 1 MHz spacing. We assume the sky is completely unpolarized. This assumption is justified because significant Faraday rotation is expected to erase any polarization response for the typical bandpass in spectral bins accessible to LuSEE-Night. We therefore simulate only the Stokes $I$ parameter.
@@
Figure 2: Simulated power beam pattern $B_{11}(\theta,\phi)$ for LuSEE-Night antenna 1 (nominally North) at 5, 15, 25, 35, and 45 MHz (left to right), derived from HFSS simulations. Projection: Orthographic, zenith-centered ($\theta=0^{\circ}$), showing the hemisphere above the lunar surface. Orientation: North (antenna 1 axis projection) is up. Scale: Power is normalized to the peak at each frequency (linear scale). Note the broad spatial response and its significant broadening towards lower frequencies. Beyond 25 MHz the antenna resonance causes the beam to bifurcate.
@@
At low radio frequencies, the system temperature is expected to be dominated by the Galactic synchrotron emission itself. We modeled the noise variance for each visibility measurement using the radiometer equation appropriate for correlation products [12]:
@@
For the simulations in this paper, we assume LuSEE-Night operates from the expected LuSEE-Night landing site at the lunar far side at lunar longitude 182.258 and latitude -23.814.
Our fiducial simulation covers one full lunar sidereal rotation period (approximately 27.3 Earth days), sampling the 16 real-valued visibility components every $\Delta t=7200$ s (2 hours). For this fiducial run, the instrument turntable was kept at a fixed orientation using the ULSA sky model and the HFSS-derived beams, adding Gaussian random noise with variance given by the radiometer equation. We therefore assume that static sky is the only present source of signal. In practice, there will also be time-varying contribution from the Sun and decametric emission from planets. We discuss how to incorporate such signals into our model in Section V.
@@ III Wiener filter map-making
  • $\mathbf{m}$ is the map vector describing the distribution of intensity on the sky. This vector could be either a map of sky intensity pixelized in an appropriate spherical pixelization (e.g., HEALPix [13]), or equivalently a set of spherical harmonics up to some finite $\ell_{\rm max}$ describing the sky intensity in harmonic space.
@@
This description is quite general and describes any instrument for which non-linearities in the instrument response can be neglected. In Wiener map-making, we further assume that the map $\mathbf{m}$ is drawn from a prior probability distribution with zero mean ($\left\langle\mathbf{m}\right\rangle=0$) and a known signal covariance matrix $\left\langle\mathbf{m}\mathbf{m}^{T}\right\rangle=\mathrm{S}$. This prior acts as a regularization, penalizing solutions that deviate significantly from the prior mean, particularly for modes where the data provide weak constraints. The balance between fitting the data and adhering to this prior is determined by the interplay between the noise covariance $\mathrm{N}$ and the signal covariance $\mathrm{S}$. This framework is mathematically equivalent to Gaussian Process Regression (GPR) [14], where the sky map $\mathbf{m}$ corresponds to the latent function to be inferred, and the signal covariance matrix $\mathrm{S}$ plays the role of the GPR kernel.
@@
Equivalently, the weighting matrix can also be written as (Method 4 of [8]):
@@ IV Results
The signal prior covariance matrix $\mathrm{S}$ was assumed diagonal in harmonic space, with $\mathrm{S}_{\ell m,\ell^{\prime}m^{\prime}}=C_{\ell}^{\rm prior}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$. We used the auto-correlation of each frequency band of the ULSA models calculated at each separate frequency. The specific characteristics of S have a minor impact in the signal-dominated regime. We employed the formulation in Eq. (13) for the Wiener filter calculation, as the number of map parameters ($N_{\rm modes}\sim\ell_{\rm max}^{2}$) was smaller than the number of data points ($N_{\rm data}$), making the inversion in map space computationally cheaper. We present a series of comparisons exploring variations from our fiducial setup. These variations include the consideration of data from different lunar cycles with varying antenna rotations, different noise levels, and the impact of beam gain fluctuations.
@@ IV.1 Recovered Sky maps
Visually, the Wiener filter successfully recovers the large-scale features of the input sky map across the frequency range. The bright Galactic Center and Galactic plane are clearly visible, as well as the attenuation coming from free-free absorption [15]. The reconstruction fidelity appears reasonably good, although some differences are noticeable, particularly at lower frequencies where the beam patterns are broadest and potentially more susceptible to modeling errors. The normalized differences indicate residuals are typically at the few percent level relative to the mean sky temperature. The reconstruction also demonstrates the filter’s behavior where data is absent: a region within approximately 24 degrees of the North Celestial Pole is unobserved by the instrument and is therefore reconstructed purely from the prior.
@@ V Conclusions & Outlook
    There will be point sources on the sky, such as Fornax A, which will be unresolved by LuSEE-Night. Since we likely know the position of those sources, a better linear model would contain a small number of known point sources, whose fluxes are fitted parameters in addition to the diffuse emission modelled as a map. First work in this direction has already started [16].
@@
    An additional extension is to model multiple frequencies. We know that neighboring frequencies will be morphologically strongly correlated. The same method can be applied by reconstructing all frequencies at once by encoding this similarity into the prior covariance matrix $\mathrm{S}$. The main issue with this approach is that the size of the data vectors would increase beyond what can be dealt with using brute-force methods. This would require a significantly more sophisticated approach than the one taken in this paper, potentially using techniques like consensus optimization to manage the computational scale (e.g., [17]).

### v2

@@ Abstract
preprint: arXiv:2508.16773
@@ I Introduction
The lunar far side has been touted as the ultimate radio observatory for low-frequency (<30<30 MHz) observations of the sky. At these low frequencies, observations from Earth are limited by both the Earth’s ionosphere, which refracts and reflects radiation from the sky, and emissions from anthropogenic and meteorological sources. The lunar far side, permanently shielded from Earth’s signals and, during the lunar night, also shielded from the Sun, has long been considered an ideal location for low-frequency radio receivers. Indeed, there are several proposals for large futuristic telescopes, such as FarSide [2021arXiv210308623B], ALO-DEX [2025arXiv250403418B], LCRT [lcrt], DSL [2021RSPTA.37990566C], as well as several recent operating pathfinders, such as the receiver on Chang’e-4 [2019EPSC...13..529S], ROLSES [2021PSJ.....2...44B, 2025arXiv250309842H] on the Intuitive Machines 1 mission, and the upcoming LuSEE-Night on the Firefly’s Blue Ghost 2 mission.
LuSEE-Night is a pathfinder radio telescope slated for delivery to the lunar far side in September 2026. A preliminary instrument description can be found in [2023arXiv230110345B], with a more complete instrument design paper in preparation. Since the monopole antennas are electrically short at the lower end of the observed frequency range and have inherently low directivity, the instrument is more akin to a radio receiver than a traditional telescope. Given the available 16 correlation products, coupled with lunar rotation and rotation of the antenna turntable, however, the dataset will contain a large number of distinct measurements. Although each one of these measurements probes large patches of the sky, their number, and signal-to-noise will overconstrain the degrees of freedom available in the low-resolution map of the sky. The purpose of this paper is to investigate this problem quantitatively and to determine if LuSEE-Night’s data can be used to recover a map of the radio sky.
To this end, we employ the Wiener filter, a well-established linear method for map reconstruction. Wiener filtering provides the maximum a posteriori (MAP) map estimate assuming a Gaussian distribution for the input map pixels and noise. It has been traditionally used in Cosmic Microwave Background (CMB) map-making [1997ApJ...480L..87T] and for general data compression in cosmology (e.g., [1998ApJ...503..492S]). In this work, we utilize Wiener filtering as a deconvolution process under a minimal set of assumptions. Future work (e.g., [Li2026]) will explore more aggressive non-linear map-making algorithms that may produce higher fidelity maps, albeit at the cost of stronger, and potentially incorrect, assumptions.
@@ II.4 Simulating the data vector
For the input sky model, we use the Ultra-Low frequency Sky Model (ULSA) [2021ApJ...914..128C], which we regenerate at frequencies from 1 to 50 MHz with a 1 MHz spacing. We assume the sky is completely unpolarized. This assumption is justified because significant Faraday rotation is expected to erase any polarization response for the typical bandpass in spectral bins accessible to LuSEE-Night. We therefore simulate only the Stokes $I$ parameter.
@@
Figure 2: Simulated beam patterns derived from HFSS simulations at 5, 15, 25, 35, and 45 MHz (columns left to right). Top row: Auto-correlation power pattern $B_{11}(\theta,\phi)$ for LuSEE-Night antenna 1 (nominally North). Middle row: Magnitude of the cross-correlation beam pattern $|B_{13}(\theta,\phi)|$ for the opposite antenna pair (1-3). Bottom row: Magnitude of the cross-correlation beam pattern $|B_{12}(\theta,\phi)|$ for the perpendicular antenna pair (1-2). Projection: Orthographic, zenith-centered ($\theta=0^{\circ}$), showing the hemisphere above the lunar surface. Orientation: North (antenna 1 axis projection) is up. Scale: All beam patterns are normalized to their peak intensity at each frequency to show relative structure (0 to 1, arbitrary units). Note the broad spatial response and its significant broadening towards lower frequencies. Beyond 25 MHz the antenna resonance causes the beam to bifurcate.
@@
At low radio frequencies, the system temperature is expected to be dominated by the Galactic synchrotron emission itself. We modeled the noise variance for each visibility measurement using the radiometer equation appropriate for correlation products [2016era..book.....C]:
@@
For the simulations in this paper, we assume LuSEE-Night operates from the expected LuSEE-Night landing site at the lunar far side at lunar longitude 182.258 and latitude -23.814. Our fiducial simulation covers one full lunar sidereal rotation period (approximately 27.3 Earth days), sampling the 16 real-valued visibility components every $\Delta t=7200$ s (2 hours). This corresponds to a continuous observation over a full lunar cycle.
While the instrument is designed to operate during both lunar day and night, the primary science operations are planned for the lunar night to minimize solar interference. The choice of a full lunar cycle for the fiducial simulation is to demonstrate the full potential of the instrument, assuming it can survive and operate for that duration. We also present results for a shorter, more pessimistic, quarter-cycle observation.
For this fiducial run, the instrument turntable was kept at a fixed orientation using the ULSA sky model and the HFSS-derived beams, adding Gaussian random noise with variance given by the radiometer equation. We therefore assume that static sky is the only present source of signal. In practice, there will also be time-varying contribution from the Sun and decametric emission from planets. We discuss how to incorporate such signals into our model in Section V.
We note that the current mission baseline includes a 50% duty cycle for the spectrometer to prevent the battery from discharging below 15% capacity over the lunar night. This operational constraint effectively halves the integration time, increasing the radiometric noise by a factor of $\sqrt{2}$ (approximately 41%). However, as demonstrated in Section IV, our map reconstruction is robust to noise variance increases of up to a factor of 10. Therefore, this modest increase in noise is not expected to significantly affect the fidelity of the reconstructed maps.
The HFSS simulations include a model of the lunar regolith as a “layered impedance boundary” consisting of an eight-layer dielectric slab. This model mimics the changing properties of the soil with depth, based on a re-analysis of Apollo 17 Surface Electrical Properties (SEP) data [Grimm2018]. It assumes a real dielectric constant increasing with depth (from $\sim 3.7$ at the surface to $\sim 6.1$ at 100 m) and a constant loss tangent of 0.01. The simulations explicitly account for the reflection from the lunar surface, including ground coupling and interference effects [Rotermund2026].
@@ III Wiener filter map-making
  • $\mathbf{m}$ is the map vector describing the distribution of intensity on the sky. This vector could be either a map of sky intensity pixelized in an appropriate spherical pixelization (e.g., HEALPix [2005ApJ...622..759G]), or equivalently a set of spherical harmonics up to some finite $\ell_{\rm max}$ describing the sky intensity in harmonic space.
@@
This description is quite general and describes any instrument for which non-linearities in the instrument response can be neglected. In Wiener map-making, we further assume that the map $\mathbf{m}$ is drawn from a prior probability distribution with zero mean ($\left\langle\mathbf{m}\right\rangle=0$) and a known signal covariance matrix $\left\langle\mathbf{m}\mathbf{m}^{T}\right\rangle=\mathrm{S}$. This prior acts as a regularization, penalizing solutions that deviate significantly from the prior mean, particularly for modes where the data provide weak constraints. The balance between fitting the data and adhering to this prior is determined by the interplay between the noise covariance $\mathrm{N}$ and the signal covariance $\mathrm{S}$. This framework is mathematically equivalent to Gaussian Process Regression (GPR) [2006gpml.book.....R], where the sky map $\mathbf{m}$ corresponds to the latent function to be inferred, and the signal covariance matrix $\mathrm{S}$ plays the role of the GPR kernel.
@@
Equivalently, the weighting matrix can also be written as (Method 4 of [1997ApJ...480L..87T]):
@@
It is important to note that while the true sky intensity is a positive quantity, the Wiener filter is a linear operation and does not enforce this constraint. This linearity ensures that the estimator remains unbiased; enforcing positivity in noise-dominated regions would rectify the zero-mean noise, introducing a bias. The resulting maps can therefore have negative values, particularly in regions or on scales where the noise is large and the solution is dominated by the zero-mean prior. Consequently, the recovery of a positive sky signal serves as a validation that the data is signal-dominated, while regions with statistically significant negative flux would indicate systematic errors. More advanced non-linear methods would be required to enforce positivity, which is beyond the scope of this paper.
@@ IV Results
The signal prior covariance matrix $\mathrm{S}$ was assumed diagonal in harmonic space, with $\mathrm{S}_{\ell m,\ell^{\prime}m^{\prime}}=C_{\ell}^{\rm prior}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$. We used the angular power spectrum of the ULSA sky model as the prior power spectrum $C_{\ell}^{\rm prior}$. We clarify that this power spectrum is computed directly from the full ULSA sky model without subtracting its mean, resulting in a non-zero power at $\ell=0$, i.e., $C_{0}\neq 0$. This $C_{0}$ is then used as the variance for the Gaussian prior on the $a_{00}$ mode, which, like all $a_{\ell m}$ priors, is centered at zero. This approach allows the Wiener filter to robustly reconstruct the sky mean when the variance $C_{0}$ is sufficiently large. The specific characteristics of S have a minor impact in the signal-dominated regime. We employed the formulation in Eq. (13) for the Wiener filter calculation, as the number of map parameters ($N_{\rm modes}\sim\ell_{\rm max}^{2}$) was smaller than the number of data points ($N_{\rm data}$), making the inversion in map space computationally cheaper. We present a series of comparisons exploring variations from our fiducial setup. These variations include the consideration of data from different lunar cycles with varying antenna rotations, different noise levels, and the impact of beam gain fluctuations.
@@ IV.1 Recovered Sky maps
Visually, the Wiener filter successfully recovers the large-scale features of the input sky map across the frequency range. The bright Galactic Center and Galactic plane are clearly visible, as well as the attenuation coming from free-free absorption [2022A&A...668A.127P]. The reconstruction fidelity appears reasonably good, although some differences are noticeable, particularly at lower frequencies where the beam patterns are broadest and potentially more susceptible to modeling errors. The normalized differences indicate residuals are typically at the few percent level relative to the mean sky temperature. The reconstruction also demonstrates the filter’s behavior where data is absent: a region within approximately 24 degrees of the North Celestial Pole is unobserved by the instrument and is therefore reconstructed purely from the prior.
@@ V Conclusions & Outlook
    There will be point sources on the sky, such as Fornax A, which will be unresolved by LuSEE-Night. Since we likely know the position of those sources, a better linear model would contain a small number of known point sources, whose fluxes are fitted parameters in addition to the diffuse emission modelled as a map. First work in this direction has already started [Fahs2026].
@@
    An additional extension is to model multiple frequencies. We know that neighboring frequencies will be morphologically strongly correlated. The same method can be applied by reconstructing all frequencies at once by encoding this similarity into the prior covariance matrix $\mathrm{S}$. The main issue with this approach is that the size of the data vectors would increase beyond what can be dealt with using brute-force methods. This would require a significantly more sophisticated approach than the one taken in this paper, potentially using techniques like consensus optimization to manage the computational scale (e.g., [2016arXiv160509219Y]).